OpenStudy (anonymous):

solve each equation x^5-8x^3+16x=o. factor by grouping?

OpenStudy (cwrw238):

first take out the x: x(x^4 - 8x^2 + 16) = 0 so one root is x = 0 now can we factor the part in the parentheses?

OpenStudy (cwrw238):

x^4 - 8x^2 + 16 = 0

OpenStudy (anonymous):

i first multiplied 1 and 16 and found the factors that multiplied to get 16 and add to -5. am i wrong?

OpenStudy (anonymous):

-4 and -4 and then i got stuck

OpenStudy (cwrw238):

- 4 * -4 = 16 and -4x^2 - 4x^2 = -8x^2 so factors are (x^2 - 4)(x^2 - 4) = 0

OpenStudy (cwrw238):

can you continue?

OpenStudy (anonymous):

where did the x^2 come from

OpenStudy (cwrw238):

x^2 * x^2 = x^4

OpenStudy (anonymous):

wouldn't i take -4 and -4 and plug it into the equation

OpenStudy (anonymous):

oh you took the x out first

OpenStudy (cwrw238):

no the next step is to solve x^2 - 4 = 0

OpenStudy (anonymous):

i dont normally do that first

OpenStudy (anonymous):

i dont normally do it that way

OpenStudy (cwrw238):

taking it out simplifies the solution

OpenStudy (anonymous):

oh

OpenStudy (cwrw238):

always look out for this - the smaller the powers the easier it is to solve

OpenStudy (cwrw238):

x^2 - 4 = 0 x^2 = 4 so x = ?

OpenStudy (anonymous):

so after taking the x out the equation out the simplified equation is x^4-8x^2=16=0

OpenStudy (anonymous):

and then whats next

OpenStudy (cwrw238):

i've already done whats next you factor to get (x^2 - 4)(x^2 - 4) = 0 so (x^2 - 4)= 0 x^2 = 4 so whats the value of x?

OpenStudy (anonymous):

2

OpenStudy (cwrw238):

2 is right but also -2 note there are 2 x^2 - 4 = 0 so we have roots of multiplicity 2 ( double roots) so our final solution set is { 0 , +/- 2, +/- 2} a total of 5 real roots

OpenStudy (cwrw238):

* 2 (x^2 - 4)

OpenStudy (anonymous):

so the answer is x^2=4; 0, +/-2, +/-2; and r real roots?

OpenStudy (anonymous):

5*

OpenStudy (cwrw238):

the answer is {0, +/- 2, +/- 2}

OpenStudy (anonymous):

thanks

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

could you give me an example of how to figure out the real zeros in a problem?

OpenStudy (anonymous):

positive negative and imaginary

OpenStudy (cwrw238):

sorry gotta go - its very late in uk

OpenStudy (anonymous):

okay im sorry thank you for your help